OpenStudy (anonymous):

Find an equation for the line: through (-5,4) and parallel to y=3/4x+3

jimthompson5910 (jim_thompson5910):

What's the slope of y = 3/4x+ 3?

OpenStudy (anonymous):

3/4 lol

jimthompson5910 (jim_thompson5910):

i'm glad you know surdawi, but I was asking katieecrowleyy

OpenStudy (anonymous):

dont think she knows

jimthompson5910 (jim_thompson5910):

Any line parallel to y=3/4x+3 will also have a slope of 3/4

jimthompson5910 (jim_thompson5910):

So plug m = 3/4, x = -5 and y = 4 into y = mx+b and solve for b